OpenStudy (anonymous):

Which of the following is equivalent to log (base 72) 9 ? log9 / log72 log72 / log9 log 8 log 1/8

ganeshie8 (ganeshie8):

use the log property :- \(\large \log_b a = \frac{\log a}{\log b}\)